LEM Deacons (Fall Term 1991)

Date Occasion  Color Preacher Service  Hymnal Deacons
10/2  St Francis  Green Scott  Lutheran  1982  SU: Brian Sliker

RC: Scott Ramming PP: John Keen

10/9  Green Susan  McIver  Orange  SU: Cindy Carr

RC: Matt Domsch PP: Kathleen Wage

10/16  Green Susan  BCP  LBW  SU: Eric Zylstra

RC: Kate Curtis PP: Jane Marston

10/23 Sharon Powell  Green Scott  Lutheran  1982  SU: Hotasi Nababan

of Lutheran RC: Brian Sliker Vol. Corp. PP: Lisa Tucker

10/30 Reformation  Red  Scott  Episc-Inc  LBW  SU: Scott Ramming

RC: John Keen PP: Matt Domsch

11/6  All Saints  White Susan  Lutheran  1982  SU: Cindy Carr

RC: Kathleen Wage PP: Eric Zylstra

11/13 Martin of Tours White Scott  BCP  LBW  SU: Kate Curtis

RC: Hotasi Nababan PP: Lisa Tucker

11/20 Oxfam  Green Susan  Lutheran  Episc  SU: Brian Sliker

RC: Jane Marston PP: Scott Ramming

11/27 Thanksgiving  Green —  Table Mass Orange  SU: Eric Zylstra

RC: Matt Domsch PP: Kathleen Wage

12/4  Advent 1  Purple Scott  Episc-Inc  1982  SU: Kate Curtis

Blue RC: Cindy Carr PP: John Keen

12/11 Advent 2  Purple Susan  Lutheran  LBW  SU: Scott Ramming

Blue RC: Hotasi Nababan PP: Brian Sliker

12/18 Advent 3  Purple Scott  BCP  1982  SU: Eric Zylstra

Blue RC: Kathleen Wage PP: Lisa Tucker